OpenStudy (anonymous):

simplify the expression. (3x^3/4)(8x^1/2) @misty1212

OpenStudy (misty1212):

is this a multiplication?

OpenStudy (anonymous):

yes

OpenStudy (misty1212):

\[\huge 3x^{\frac{3}{4}}\times 8x^{\frac{1}{2}}\]?

OpenStudy (misty1212):

\[3\times 8=24\] is a start then add up the exponents \[\frac{3}{4}+\frac{1}{2}=\frac{3}{4}+\frac{2}{5}=\frac{5}{4}\]

OpenStudy (misty1212):

final answer \[\huge 24x^{\frac{5}{4}}\]